# On the mechanics of inhaled bronchial transmission of pathogenic microdroplets generated from the upper respiratory tract, with implications for infection onset

### v1

@@ II.1.3 Numerical experiments for particulate transport
Here $u_{pi}$ represents the particulate velocity, $\rho_{p}$ is the material density of the particulates, $d$ represents the particulate diameter, $Re_{p}$ is the particulate Reynolds number, $g_{i}$ signifies the gravitational acceleration in the $i$ direction, $C_{D}$ is the drag coefficient, and $F_{i}$ is any other additional body forces per unit particulate mass, e.g., the Saffman lift force exerted by a typical flow-shear field on small particulates transverse to the airflow direction. In this context, note that the solution scheme considered the particulates to be large enough to ignore any Brownian motion effects on their dynamics.
To evaluate the drag term (i.e., the co-factor on the first term) on the right hand side of equation 10, the quantities $Re_{p}$ and $C_{D}$ are respectively computed as
$Re_{p}$ $=\frac{\rho_{p}d\left|u_{i}-u_{pi}\right|}{\mu}$ $C_{D}$ $=a_{1}+\frac{a_{2}}{Re_{p}}+\frac{a_{3}}{{Re_{p}}^{2}},$ (11)
where $\mu$ is the molecular viscosity of the ambient fluid (i.e., air), while $a_{1}$, $a_{2}$, and $a_{3}$ are functions of $Re_{p}$ determined based on the spherical drag law [28]. Subsequently, the particulate trajectories are derived from their spatiotemporal locations, $x_{i}\left(t\right)$, obtained by numerical integration of the following velocity vector equation:
@@
Intra-airway aerial tracking of a particulate pathline was stopped once it entered the mesh element layer adjacent to the enclosing walls of the respiratory cavity. The particulates escaping through the bronchiolar outlets were also recorded, and represented the particulates penetrating to the respiratory bronchioles and deep lungs. The numerical experiments in this study tested particulates of diameters $\mathrm{1-30~{}\mu m}$ (with increments of 1 $\mathrm{\mu m}$). In total, $\mathcal{N}$ = 1622 particulates of each size were tracked; $\mathcal{N}$ being the number of surface elements in the mesh layer mapping the inlet face (see Fig. 1). Also, considering that saliva-mixed mucus is 99.5% water [29], the material density of the particulates was assumed to be $\rho_{p}$ = 1.00175 g/ml, approximated as the weighted average (see equation 13 below) between 99.5% water and the residual pathogenic and biological non-volatile compounds with a representative density of $\rho_{nv}=$ 1.35 g/ml; e.g., for protein [30], the density being independent of the nature of the protein and particularly independent of its molecular weight. The weighted average calculation for the simulated particulate density is as follows:
@@
Figure 3: (a) Laryngotracheal region upwind from the primary bronchi used to develop a simulation-informed reduced-order analytical model (S-ROAM) for bronchial transmission. AA marks the distal larynx upwind face and BB marks the downwind face of the distal tracheal cavity; see labels in Fig. 1d. The cross-sectional areas of the cavity at AA and BB are 108.61 mm2 and 123.03 mm2, respectively. The corresponding hydraulic diameters are 11.03 mm and 12.20 mm, respectively. The linear spatial distance between the two faces is approximately 115.84 mm, at angle $\theta_{1}$ = 23.77 to the vertical (direction of gravity in the numerical simulations). Panel (b) highlights the location and orientation of the planar cross-section shown in panel (c) and also previously as the location of the mesh visuals in Fig. 2. The vortex strengths and positions for the S-ROAM are extracted from the simulated data mapping this two-dimensional plane; the plane cuts through the entire cavity of panel (a). $\theta_{2}$ = 33.94 is the assumed angle in S-ROAM, at which microdroplets enter the AA inlet face and is governed by the anatomical shape of the cavity upwind from AA (marked by the dashed black traces; also see Fig. 1).
@@
II.2 Analytical model for inhaled airflow and particulate transport
@@
To verify the particulate transport trends to the bronchial domain as derived numerically, this study invokes a simulation-informed reduced-order analytical model (S-ROAM), with a two-dimensional channel mimicking the laryngotracheal domain. Therein, the inhaled airflow was modeled in the dimensionless complex $\chi$ plane, with $\chi=\alpha+\mathrm{i}\,\beta$ and $\mathrm{i}\,\beta$ aligned with the channel’s streamwise axis; see Fig. 3. The two-dimensional channel had its streamwise length and cross-stream width based on the averaged dimensions of the anatomical tract from Fig. 3a-b. The channel walls were placed at $\alpha=0,1$. Based on the geometric inputs, the S-ROAM enforced $L/W\approx 9$, where $L$ represents the streamwise channel length and $W$ is the inlet width. For additional details on the analytical approach, see our recent paper [31].
@@
The reduced system modeled an extended vortex patch emerging in the full-scale numerical simulation as a collection of five point vortices distributed over the area of the simulated vortex core (on the channel plane, mapping the mid-sagittal section of the three-dimensional cavity; see Fig. 3b), bearing dimensionless circulations $\Gamma_{i}=\omega A/5UW$, where $\omega$ was the mean vorticity (as determined from the numerical simulations) spread over the patch area $A$ and $U$ is the characteristic streamwise air speed through the channel. From the simulated data, we had: $\omega=1750$ s-1, $A=2.65\times 10^{-5}$ m2, and $U=Q/a=2.16$ m/s, with $a=$ 115.82 mm2 being the area of the inlet face and $Q=15$ L/min $=0.00025$ m3/s being the simulated inhalation flux. The anatomical planar section from which the vorticity information was extracted is shown in Fig. 3c.
@@
with $\phi$ as the velocity potential and $\psi$ as the real-valued flow streamfunction; the asterisk denotes complex conjugation.
Subsequently, the inhaled particulate motion was analytically derived using a simplified version of the Maxey-Riley equation [35], in the two-dimensional vector form [36, 37]:
@@
Although viscous effects are neglected in the above velocity model, particle drag is included. This form of the model ignores the Faxen correction terms and the Basset-Boussinesq history force. Particulates are assumed to be entering the S-ROAM domain (inclined at an angle $\theta_{1}=23.77^{\circ}$ to the direction of gravity, per the general shape and orientation of the test anatomical cavity) with speed $U$ and at an angle of $\theta_{2}=33.94^{\circ}$ with respect to the S-ROAM’s $\alpha$-axis, mimicking the simulated motion of particles entering the region; see Fig. 3c. As representative examples, particulate trajectories were derived for two initial positions at the mid-point and near the right edge of the reduced channel’s inlet span, and for seven different particulate diameters, namely, 5, 10, 15, 20, 25, 30, and (as an extremal case) 50 $\mathrm{\mu m}$.
II.3 Connecting the fluid dynamics outcomes to virology
From the numerical experiments, this study has deduced the simulated deposition and penetration efficiencies of the particulates along the bronchial tubes (see Table 1) and evaluated the respective volumetric transmission to the bronchi. The projected net deposited (at the primary and tertiary bronchi) and penetrated (moving into the respiratory bronchioles) volumes are then multiplied with the sputum viral concentration for a specific pathogen (in this study, SARS-CoV-2) to evaluate and compare the viral load transported via inhaled aerial advection of URT-derived liquid particulates to the lower airway and deep lungs, for select test particulate sizes, namely, 1, 5, 10, and 15 $\mathrm{\mu m}$.
Figure 4: 30 representative velocity streamlines extracted from the numerically simulated inhaled airflow field.
Note that the average sputum viral concentration for SARS-CoV-2 has been reported as $\mathcal{V}$ = 7.0$\times$106 virions/ml, through count measurements of the RNA copies for the single-stranded virus present in the airway liquid samples collected from hospitalized COVID-19 patients [7].
@@ III Results
Table 1: Numerically simulated bronchial deposition and penetration data: The particulate sizes, that have been explored further in Table 2 (to formalize the viral transmission trends as a function of the microdroplet dimensions), are in bold font. Symbols: $d$ = tested aerosol (or, droplet) diameter; $\mathcal{N}$ = total number of aerosols/droplets tracked for each particulate diameter; $n_{p}$ = number of deposited particulates in the primary bronchi; $n_{t}$ = number of deposited particulates in the tertiary bronchi; $n_{d}$ = number of particulates penetrating into the respiratory bronchioles toward the deep lungs; $\eta$ = deposition (or, penetration) rate (in %) to the bronchial pathways.
Separated aerosol/droplet size
@@
III.1 Bronchial deposition and deep lung penetration as projected from numerical modeling
Figure 4 presents sample inhaled airflow velocity streamlines from the numerical simulation of inhaled airflow passing through the CT-reconstructed anatomical cavity. The simulated inhalation rate of 15 L/min to mimic relaxed breathing warranted an inlet-to-outlet pressure gradient of -9.5 Pa. Against the flow field, the test particulates were tracked with their starting locations uniformly distributed on the cross-sectional space spanning the vocal fold region of the laryngeal cavity (commensurate with their hypothesized formation sites at the URT—through break-up of mucus strata along the nasopharynx, the oropharynx, and the vocal folds).
Table 1 details the lower airway deposition and penetration data for the tested particulates bearing diameters $\mathrm{1-30~{}\mu m}$ (with increments of $\mathrm{1~{}\mu m}$). While it is expected that the particulates $\lesssim$ $\mathrm{5~{}\mu m}$ would relatively comfortably penetrate to the deep lungs (as is clearly the case per Table 1), the high transmission percentages, e.g., for even the 10- and 15-$\mathrm{\mu m}$ particulates is striking—them being 88.41% and 60.85%, respectively. In the table (for 1622 particulates tracked for each size), $n_{p}$ quantifies the number of particulates depositing in the primary bronchi, $n_{t}$ quantifies the number of particulates depositing at the tertiary bronchi, and $n_{d}$ measures the number of particulates escaping through the respiratory bronchiolar outlets of the in silico model and subsequently moving into the deeper recesses of the lungs. See Fig. 1 to review the anatomical zone nomenclature for this study and the respective locations of the different physiological regions along the airway.
@@ III.2 Viral load transmitted to the lower airway
Table 2 lists the viral load transmitted to the bronchial pathways, for the representative test particulate sizes 1, 5, 10, and 15 $\mathrm{\mu m}$. For the time scale $T$ (in hours) over which the viral load transmission is to be estimated, we have used 3 days (i.e., 72 hours) based on reported data [38] on the typical time interval that has revealed confirmed infection onset in the deep lungs subsequent to the emergence of initial symptoms along the URT. Further, applying $\delta t$ = 5 s as the average duration for a complete breathing cycle [39], the viral load, $L_{v}$, transmitted to the bronchial pathways and deep lungs can be computed as:
@@
where $\mathcal{P}\left(d_{i}\right)$ is the particle size density function that quantifies the number of microdroplets of size $d_{i}$ generated during each inhalation cycle. For simplicity, the data in Table 2 enforced $\mathcal{P}$ = 1, irrespective of the particulate size, i.e., the particulate formation rate was conservatively assumed to be 1 per breathing cycle, for each test particulate size. Finally, note that while the reported numbers in Table 2 are based on the estimates guided by equation 16, the zero viral load assessment for the 1-$\mathrm{\mu m}$ particulates implies that the corresponding transmission estimate from equation 16 resulted in a fractional number $\ll$ 1, and bears no physical significance.
@@
Separated aerosol/droplet size
@@
Number of virions ferried to the bronchial pathways
@@
Figure 5: Log-log plot for $\eta$, summing up the bronchial deposition and deep lung penetration percentages (see Table 1), for each numerically tracked microdroplet size, $d$. The simulation-derived data points are marked by the dark blue circles. The solid black line is a fitted curve of the Heaviside step function form (see equation 17).

### v2

@@ II.0.3 Numerical experiments for particulate transport
Here $u_{pi}$ represents the particulate velocity, $\rho_{p}$ is the material density of the particulates, $d$ represents the particulate diameter, $Re_{p}$ is the particulate Reynolds number, $g_{i}$ signifies the gravitational acceleration in the $i$ direction, $C_{D}$ is the drag coefficient, and $F_{i}$ represents additional body forces per unit particulate mass in the form of the Saffman lift force exerted by a typical flow-shear field on small particulates transverse to the airflow direction. In this context, note that the solution scheme considered the particulates to be large enough to ignore any Brownian motion effects on their dynamics.
To evaluate the drag component in equation 11, the quantities $Re_{p}$ and $C_{D}$ are respectively computed as: $Re_{p}=\rho_{p}d\left|u_{i}-u_{pi}\right|/\mu$ and $C_{D}=a_{1}+a_{2}/Re_{p}+a_{3}/{Re_{p}}^{2}$; where $\mu$ is the molecular viscosity of the ambient fluid (i.e., air), while $a_{1}$, $a_{2}$, and $a_{3}$ are functions of $Re_{p}$ determined based on the spherical drag law [48]. Subsequently, the particulate trajectories are derived from their spatiotemporal locations, $x_{i}\left(t\right)$, obtained by numerical integration of the following velocity vector equation:
@@
Intra-airway aerial tracking of a particulate pathline was stopped once it entered the mesh element layer adjacent to the enclosing walls of the respiratory cavity. The particulates escaping through the bronchiolar outlets were also recorded, and represented the particulates penetrating to the respiratory bronchioles and deep lungs. The numerical experiments in this study tested particulates of diameters $\mathrm{1-30~{}\mu m}$ (with increments of 1 $\mathrm{\mu m}$). In total, $\mathcal{N}$ = 1622 particulates of each size were tracked; $\mathcal{N}$ being the number of surface elements in the mesh layer mapping the inlet face (see Fig 1). Also, considering that saliva-mixed mucus is 99.5% water [49], the material density of the particulates was assumed to be $\rho_{p}$ = 1.00175 g/ml, approximated as the weighted average (see equation 13 below) between 99.5% water and the residual pathogenic and biological non-volatile compounds with a representative density of $\rho_{nv}=$ 1.35 g/ml; e.g., for protein [50], the density being independent of the nature of the protein and particularly independent of its molecular weight. The weighted average calculation for the simulated particulate density is as follows:
@@
II.0.4 Supplementary ‘diversion’: bronchial transmission trend for particulates inhaled from outside
While the focus of this study is singularly on the downwind bronchial transmission of particulates generated within the URT, it would be of scholarly interest to undertake a brief detour at this point to explore an auxiliary question (that has been nonetheless well-explored in literature): for the particulates inhaled from the external environment, which particulate sizes would be efficient at penetrating to the bronchial spaces? Here, we address this question specifically for the test geometry used in this study, through transport simulations within its non-redacted complete anatomy (Fig 1b) wherein inhaled particulates are entering the airway through both the left and right nostrils.
Comparing the cavity spaces in the complete airway geometry (Fig 1b) and the redacted model (Fig 1c-d), the volume of the former is approximately 3.48 times larger than that of the redacted system. Accordingly, following the described meshing protocol, the full airway was discretized into 2.5 million $\times$ 3.48 = 8.7 million unstructured tetrahedral elements, with four layers of pentahedral cells lining the cavity surfaces. Using this meshed geometry, particulate transport was simulated at an inhaled airflow rate of 15 L/min, following methods outlined earlier. The only modification was setting the material density of the inhaled particulates to 1.3 g/ml, consistent with typical values for environmentally dehydrated respiratory ejecta [49, 25], which constitutionally form the pathogen-bearing particulates inhaled from external air.
Fig 3 demonstrates the bronchial transmission trend for particulates inhaled from outside. Therein, panel (a) shows the cumulative transmission percentage, $\eta$, at the primary, secondary, and tertiary bronchi, together with the penetration rate (in %) through the geometry outlets that serve as entry to the respiratory bronchioles. Panel (b) solely presents $\eta_{d}$, the penetration rate through the bronchiolar outlets, indicating the proportion of inhaled particulates reaching the deeper lung regions. Considering select particulate sizes, $\eta$ starts at 99.36% for particulates of diameter $d=1~{}\mathrm{\mu m}$, decreases to 96.86% for $d=5~{}\mathrm{\mu m}$, to 31.46% for $d=10~{}\mathrm{\mu m}$, and approaches 0% for $d=15~{}\mathrm{\mu m}$ and larger. Remarkably though, as we will find later (in the results), for particulates generated within the URT, the quantity $\eta$ shoots up significantly, approaching 90% for $d=10~{}\mathrm{\mu m}$ and steadily sustaining at $>$ 60% even for $d=15~{}\mathrm{\mu m}$.
Figure 3: Transmission trend for particulates inhaled from outside and navigating the complete anterior airway. (a) $\eta$ represents the cumulative deposition percentages at the primary, secondary, and tertiary bronchi, together with the penetration rate through the bronchiolar outlets into the deeper lung regions; (b) $\eta_{d}$ isolates the penetration rate through the bronchiolar outlets with the tracked particulates moving into the deeper lung regions. The reader may find it insightful to compare the trend reported here with Fig 6a-b on the navigation trend for particulates generated within the URT (and still airborne at the larynx). In the latter scenario, the larger particulates exhibit much greater efficiency at penetrating to the lower airway.
@@ II.0.4 Supplementary ‘diversion’: bronchial transmission trend for particulates inhaled from outside
II.1 A simplified analytical model for inhaled transport through the laryngotracheal space
@@
To verify the particulate transport trends to the bronchial domain as derived numerically, let us invoke a simulation-informed reduced-order analytical model (S-ROAM), with a two-dimensional channel mimicking the laryngotracheal domain. Therein, the inhaled airflow is modeled in the dimensionless complex $\chi$ plane, with $\chi=\alpha+\mathrm{i}\,\beta$ and $\mathrm{i}\,\beta$ aligned with the channel’s streamwise axis; see Fig 4. The two-dimensional channel has its streamwise length and cross-stream width based on the averaged dimensions of the anatomical tract from Fig 4a-b. The channel walls are placed at $\alpha=0,1$. Based on the geometric inputs, the S-ROAM enforces $L/W\approx 9$, where $L$ represents the streamwise channel length and $W$ is the inlet width. For a detailed exposition of this classical modeling approach for complex respiratory systems, see our recent preprint [51].
Figure 4: Defining the analytical domain, with representative numerical visuals. (a) Laryngotracheal region upwind from the primary bronchi used to develop a simulation-informed reduced-order analytical model (S-ROAM) for bronchial transmission. AA marks the larynx upwind face and BB marks the downwind face of the distal tracheal cavity; see labels in Fig. 1d. The cross-sectional areas of the cavity at AA and BB are 108.61 mm2 and 123.03 mm2, respectively. The corresponding hydraulic diameters are 11.03 mm and 12.20 mm, respectively. The linear spatial distance between the two faces is approximately 115.84 mm, at angle $\theta_{1}$ = 23.77 to the vertical (direction of gravity in the numerical simulations). Panel (b) highlights the location and orientation of the planar cross-section shown in panel (c) and also previously as the location of the mesh visuals in Fig 2. The vortex strengths and positions for the S-ROAM are extracted from the simulated data mapping this two-dimensional plane; the plane cuts through the entire cavity of panel (a). $\theta_{2}$ = 33.94 is the assumed angle in S-ROAM, at which microdroplets enter the AA inlet face and is governed by the anatomical shape of the cavity upwind from AA (marked by the dashed black traces; also see Fig 1). (d) 30 randomly selected representative velocity streamlines extracted from the numerically simulated inhaled airflow field. (e) Simulated trajectory of a representative 10-$\mathrm{\mu m}$ particulate. The geometric centroid of the inlet face, marked by the solid circle, was chosen as the particulate’s position at tracking time $t=0$.
@@
The reduced system modeled the core of an extended vortex patch emerging in the full-scale numerical simulation as a random collection of five point vortices distributed over the area of the simulated vortex core (on the channel plane, mapping the mid-sagittal section of the three-dimensional cavity; see Fig 4b), bearing dimensionless circulations $\Gamma_{i}=\omega A/5UW$, where $\omega$ was the mean vorticity (as determined from the numerical simulations) spread over the patch area $A$ and $U$ is the characteristic streamwise air speed through the channel. From the simulated data, we had: $\omega\approx 1750$ s-1, $A\approx 2.65\times 10^{-5}$ m2, and $U=Q/a=2.16$ m/s, with $a=$ 115.82 mm2 being the area of the inlet face and $Q=15$ L/min $=0.00025$ m3/s being the simulated inhalation flux. Based on the angular orientation of the modeled core in the simulated vorticity field, the model point vortices were placed on a line that subtended $\approx$ 2.2 in counter-clockwise sense with the vertically downward direction. The anatomical planar section from which the vorticity information was extracted is shown in Fig 4c.
@@
with $\phi$ as the velocity potential and $\psi$ as the real-valued flow streamfunction; the asterisk denotes complex conjugation. Subsequently, the inhaled particulate motion was analytically derived using a simplified version of the Maxey-Riley equation [55, 56], in the two-dimensional vector form [57, 58]:
@@
Although viscous effects are neglected in the above velocity model, particle drag is included. This form of the model ignores the Faxen correction terms and the Basset-Boussinesq history force. Particulates are assumed to be entering the S-ROAM domain (inclined at an angle $\theta_{1}=23.77^{\circ}$ to the direction of gravity, per the general shape and orientation of the test anatomical cavity) with speed $U$ and at an angle of $\theta_{2}=33.94^{\circ}$ with respect to the S-ROAM’s $\alpha$-axis, mimicking the simulated motion of particles entering the region; see Fig 4c. As representative examples, particulate trajectories were derived for two initial positions at the mid-point and near the right edge of the reduced channel’s inlet span, and for seven different particulate diameters, namely 5, 10, 15, 20, 25, 30, and (as an extremal case) 50 $\mathrm{\mu m}$.
II.2 Connecting the fluid dynamics outcomes to pathogen-specific virological data
From the numerical experiments, this study has deduced the simulated deposition and penetration efficiencies of the URT-derived particulates along the bronchial tubes (see Table 1) and evaluated the respective volumetric transmission to the bronchi. The projected net deposited (at the primary, secondary, and tertiary bronchi) and penetrated (moving into the respiratory bronchioles) volumes are then multiplied with the sputum viral concentration for a specific pathogen (in this study, SARS-CoV-2) to evaluate and compare the viral load transported via inhaled aerial advection of URT particulates to the lower airway and deep lungs, for select test particulate sizes, namely, 1, 5, 10, and 15 $\mathrm{\mu m}$.
Note that the average sputum viral concentration for SARS-CoV-2 has been reported as $\mathcal{V}$ = 7.0$\times$106 virions/ml, through count measurements of the RNA copies for the single-stranded virus present in the airway liquid samples collected from hospitalized COVID-19 patients; see [23].
@@ III Results
Table 1: Numerical transmission trend: The numerically simulated bronchial deposition and penetration data are detailed herein. The particulate sizes, that have been explored further in Table 2 (to formalize the viral transmission trends as a function of the microdroplet dimensions), are in bold font. Symbols: $d$ = tested aerosol (or, droplet) diameter; $\mathcal{N}$ = total number of aerosols (or, droplets) tracked for each particulate diameter; $n_{p}$ = number of deposited particulates in the primary and secondary bronchi; $n_{t}$ = number of deposited particulates in the tertiary bronchi; $n_{d}$ = number of particulates penetrating into the respiratory bronchioles toward the deep lungs; $\eta$ = cumulative deposition (or, penetration) rate (in %) to the bronchial pathways and computed as $100\times\left(\sum_{j}n_{j}\right)/\mathcal{N}$, with $j\in\{p,t,d\}$.
Separated particulate size
@@
III.1 Bronchial deposition and lower airway penetration as projected from numerical modeling
Fig 4d presents sample inhaled airflow velocity streamlines from the numerical simulation of inhaled airflow. The simulated flux of 15 L/min warranted an inlet-to-outlet pressure gradient of -9.5 Pa. Against the flow field, the test particulates were tracked with their starting locations uniformly distributed on the cross-sectional space spanning the vocal fold region of the laryngeal cavity (commensurate with their hypothesized formation sites at the URT—through breakup of mucus strata along the nasopharynx, the oropharynx, and the vocal folds). A representative particulate trajectory for $d=10~{}\mathrm{\mu m}$ has been additionally shown in Fig 4e. At tracking time $t=0$, the particulate was assumed to be positioned at the geometric centroid of the inlet face. Eventually, the sample particulate penetrates through the outlet to respiratory bronchioles, thereby moving into the deeper lung recesses.
Fig 5a-c and Table 1 detail the lower airway deposition and penetration data for the tested particulates bearing diameters $\mathrm{1-30~{}\mu m}$ (with increments of $\mathrm{1~{}\mu m}$). While it is expected that the particulates $\lesssim$ $\mathrm{5~{}\mu m}$ would comfortably penetrate to the deeper regions of lungs (as is clearly the case per Table 1; see the top rows), the high transmission percentages, e.g., for even the 10- and 15-$\mathrm{\mu m}$ particulates is striking—them being 88.04% and 61.34%, respectively.
Figure 5: Numerically simulated transmission trend for URT-derived particulates navigating through the redacted test geometry. (a) $\eta$ represents the cumulative deposition percentages at the primary, secondary, and tertiary bronchi, as well as the penetration rate through the bronchiolar outlets into the deeper lung regions; (b) $\eta_{d}$ isolates the penetration rate through the bronchiolar outlets; (c) The solid red line is a fitted curve of the Heaviside step function form (see equation 17), with the solid black circles showing simulation-derived $\eta$ values (reported in Table 1).
@@ III.2 Viral load transmitted to the lower airway
Table 2 lists the viral load transmitted to the bronchial pathways, for the representative test particulate sizes 1, 5, 10, and 15 $\mathrm{\mu m}$. For the time scale $T$ (in days) over which the viral load transmission is to be estimated, we have used 3 days based on reported data [59] on the typical time interval that has revealed confirmed infection onset in the deep lungs subsequent to the emergence of initial symptoms along the URT. Further, applying $\delta t$ = 5 s as the average duration for a complete breathing cycle [60], the viral load, $L_{v}$, transmitted downwind can be computed as:
@@
where $\mathcal{P}\left(d_{i}\right)$ is the particle size density function that quantifies the number of microdroplets of size $d_{i}$ generated during each inhalation cycle. For simplicity, the data in Table 2 enforced $\mathcal{P}$ = 1, irrespective of the particulate size, i.e., the particulate formation rate was conservatively assumed to be 1 per breathing cycle, for each test particulate size. Also, while the reported numbers in Table 2 are based on the estimates guided by equation 16, the zero viral load assessment for the 1-$\mathrm{\mu m}$ particulates implies that the corresponding transmission estimate from equation 16 resulted in a fractional number $\ll$ 1, and bears no physical significance. To note here additionally, for the assumed $\mathcal{P}$, the quantity $L_{v}$ exceeded 300 for $d\in[13,22]~{}\mathrm{\mu m}$. See the conclusion section later for the related translational relevance in infection mechanics.
@@
Separated particulate size
@@
Number of virions ferried
to the bronchial pathways